# Neutron-nucleus dynamics simulations for quantum computers

### 22 Feb 2024 v1

@@ VI.3 Description of the quantum simulations
The ansatz choice for the different simulations is given in Sec. IV.2. For the n+C systems with an exponential potential, we perform simulations with $N=8$ and $N=16$, with $K=3$ (7-diagonal matrix). For the n+$\alpha$ systems, we use $N=8$, with $K=1$ and $K=2$. The type of the simulations are as follows:
  • Exact diagonalization: We diagonalize the matrix exactly (using classical methods) and find the minimum eigenvalue, which is used to validate the quantum simulation outcomes.
@@
  • Noise-resilient estimation: We use the final parameters $\bm{\theta}$ of the noisy simulation from the noisy simulation and calculate the expectation value of the Hamiltonian on a classical machine to find a noiseless estimate for the minimum energy, which we refer to as the noise-resilient (NR) value. The idea of noise-resilient training was first introduced in [22]. We find that the algorithm exhibits noise resilience; i.e., the NR value is more accurate than the noisy simulation outcome. In other words, training is still possible in a noisy scenario.
@@
For the n+C system, we take as input $N$, $K$, $\hbar\omega$, $V_{0}$, and $c$. For the n+$\alpha$ system, we take as input $N$, $K$, $\hbar\omega$, and the list of coefficients $\{v_{k}\}_{k=0}^{K}$.
@@ VI.4 Quantum simulations for neutron-Carbon dynamics
In this section we present the results of the different simulations for n-Carbon dynamics. The specific details, including the number of iterations, type of gradient estimator, and number of shots, can be found in Appendix C. We provide simulations that, for the first time, show the efficacy of the Gray code for a Hamiltonian matrix beyond the tridiagonal case, that is, for $K=3$ (bandwidth of 7). With the Gray code, we can utilize only three and four qubits to simulate model spaces of $N=8$ and 16 basis states, respectively. This provides acceptable results that closely agree with the energy for the $K=3$ approximation, which lies near the exact theoretical energy as discussed above (see Fig. 9 for $N=8$ and $N=16$). The quantum simulations for n$+^{10}$, n$+^{12}$, and n$+^{14}$ are shown in Figs. 11-13, and the case of n$+^{14}$ is compared to the one-hot encoding in Fig. 14.
@@
Encoding $E_{\rm th}$ $N$ $E_{{\rm th}}^{K=3}$ $E^{K=3}_{\rm NL}$ $E^{K=3}_{\rm shot}$ $E^{K=3}_{\rm noisy}$ $E^{K=3}_{\rm NR}$
[MeV] [MeV] [MeV] [MeV] [MeV]
n+${}^{10}$ Gray $-6.78$ $8$ $-6.5364$ $-6.5364\pm 2\times 10^{-10}$ $-6.5305\pm 0.0005$ $3.5\pm 0.5$ $-5.8567$
$16$ $-6.7346$ $-6.668\pm 0.007$ $-6.62\pm 0.02$ $12.6\pm 0.7$ $-6.009$
n+${}^{12}$ Gray $-1.86$ $8$ $-1.18495$ $-1.184993\pm 1\times 10^{-6}$ $-1.1701\pm 0.0006$ $8.7\pm 0.5$ $-0.5264$
$16$ $-1.70020$ $-1.6973\pm 0.0001$ $-1.599\pm 0.003$ $13.9\pm 0.7$ $-0.7637$
n+${}^{14}$ Gray $-1.22$ $8$ $-0.49963$ $-0.4996\pm 6\times 10^{-10}$ $-0.4966\pm 0.0005$ $8.4\pm 0.3$ $0.1103$
$16$ $-1.0070$ $-0.9860\pm 0.0005$ $-0.982\pm 0.002$ $15.2\pm 0.5$ $0.5942$
One-hot $8$ $-0.49963$ $-0.4987\pm 0.0001$ $-0.49\pm 0.03$ $9.0\pm 0.6$ $0.7246$
Table 4: Simulation energy for the lowest $\frac{1}{2}^{+}$ state for each neutron-Carbon system, for the $K=3$ approximation with $N=8$ or $N=16$, corresponding to Figs. 11-13: the true value $E_{\rm th}^{K=3}$; $E^{K=3}_{\rm NL}$ from the noiseless simulations, $E^{K=3}_{\rm shot}$ form the shot-noise simulations, and $E^{K=3}_{\rm noisy}$ from the noisy simulations, reported as the mean and standard deviation of the last 100 iterations; and the NR value $E^{K=3}_{\rm NR}$. The theoretical eigenenergy $E_{\rm th}$ of the exact potential is also shown (cf. Table 3).
In particular, in all cases the noiseless and shot-noise simulations with the Gray code yield results with very small errors, mostly, $10^{-10}$-$10^{-3}$. Importantly, they reproduce the true value within 00-$2$%, with the only exception being the $6$-deviation in $E^{K=3}_{\rm shot}$ for ${}^{12}$ and $N=16$.
A very significant result is that the noisy simulations, which are expected to simulate runs on current NISQ processors, provide very reasonable energy estimates through the use of the NR value. In particular, we find that the algorithm exhibits noise resilience; i.e., the NR value $E^{K=3}_{\rm NR}$ is more accurate than the noisy simulation outcome and training of the wavefunctions parameters is still possible in a noisy scenario. Furthermore, for these n-C systems, we show that the $E^{K=3}_{\rm NR}$ energies differ from the corresponding true value only by $600$-$950$ keV, which surpasses the accuracy of many nuclear models. The only exception is the four-qubit simulation for n$+^{14}$ (Fig. 13), which predicts an unbound $\frac{1}{2}^{+}$ state at $0.59$ MeV for the weakly bound state at $-1.0$ MeV. While this is still a reasonable estimate, weakly bound states require larger model spaces to achieve convergence, as discussed above, which implies the need for larger number of qubits. This means that special care needs to be taken in the quantum simulations of such systems, including n$+^{16}$ and n$+^{18}$, when performed on NISQ devices. One way to improve this is to try different ansatz structures, or a larger model space, which remains to be shown in future work.
Nonetheless, even in the case of weakly bound states, the Gray code with three qubits is superior to the one-hot encoding with eight qubits, as illustrated in Fig. 14 and Table 4. Indeed, compared to the Gray-code case, the one-hot simulations are much slower and show worse performance in the presence of noise. In particular, the one-hot simulation yields larger errors [e.g., by six (two) orders of magnitude for the noiseless (shot noise) simulations], as well as a deviation for $E^{K=3}_{\rm NR}$ from the true value that is twice as large as the corresponding Gray-code estimate. In addition, as shown in Figs. 2 and 3 as well as in Tables 1 and 2, the one-hot ansatz utilizes 13 two-qubit gates compared to only 8 two-qubit (CNOT) gates in the Gray encoding ansatz with $L=4$ used in our simulations. This suggests that the Gray encoding and the use of fewer qubits are indeed highly advantageous for nuclear problems that achieve convergence in larger model spaces, with the case of weakly bound systems presented here being an illustrative example.
Figure 14: Simulation using the one-hot encoding for the energy of the lowest $\frac{1}{2}^{+}$ state for n$+^{14}$ modeled by the exponential potential Eq. (13) with $N=8$ ($n=8$ qubits) and $K=3$. The inset plot is the last $200$ iterations.
Figure 15: Simulation for the n-$\alpha$ ab initio potential for $\hbar\omega=12$ MeV Eq. (VI.2) with $N=8$, and (a) $K=1$ and (b) $K=2$. The inset plot is the last 200 iterations.
VI.5 Quantum simulations for n$+\alpha$ using ab initio optical potentials
In this section, we present the results of the simulations for the n+$\alpha$ dynamics. The specific details, including the number of iterations, type of gradient estimator, and number of shots, can be found in Appendix D. The quantum simulations for n$+\alpha$ are carried out for model spaces of $N=8$ (Fig. 15) and $N=16$ (Fig. 16) for $\hbar\omega=12$ MeV and of $N=8$ (Fig. 17) for $\hbar\omega=16$ MeV. We show the cases of $K=1$ for a potential of $\mathcal{O}(r^{4})$ and $K=2$ for a potential of $\mathcal{O}(r^{6}$).
We summarize the energy estimates from the various quantum simulations in Table 5. The energies $E_{\rm th}^{K}$ from the exact diagonalization for $K=1$ and $2$ converge within the first four to five digits with the increasing model space from $N=8$ to $N=16$, as illustrated in Table 5 for $\hbar\omega=12$ MeV (for $\hbar\omega=16$ MeV, $E_{\rm th}^{K=1}=-20.77$ MeV and $E_{\rm th}^{K=2}=-18.95$ MeV for $N=16$). Hence, the final estimates are based on the $N=8$ results, where the uncertainties are estimated for a 14% variation of the $\hbar\omega$ values ($\hbar\omega=12$-$16$ MeV), the same range that yields a total cross section for the neutron scattering on ${}^{4}$e that is converged and in agreement with experiment, as discussed in Ref. [23]. In addition, the $K=1$ and $K=2$ approximation energies are very close to the corresponding $E_{0}$ eigenenergy of the original non-local ab initio optical potential. In fact, across the $\hbar\omega$ range, the energy of the lowest $\frac{1}{2}^{+}$ orbit is estimated at $-19.3(1.5)$ MeV for $K=1$ and at $-17.8(1.2)$ MeV for $K=2$, both of which agree within the uncertainties with the non-local estimate of $-19.8(1.0)$ reported above. Exactly the same energy estimates are obtained by the noiseless and shot simulations ($-19.3\pm 1.5$ MeV for $K=1$ and at $-17.8\pm 1.2$ MeV for $K=2$). We note that the errors that stem from these simulations (of order of a few keV) are inconsequential compared to the ones associated with the $\hbar\omega$ variation ($\sim 1-1.5$ MeV). Importantly, while the noisy simulations yield unacceptably large energies, the corresponding NR energies agree with the shot-noise outcomes within $1\sigma$: $-18.6(1.7)$ MeV for $K=1$ and $-17.1(1.2)$ MeV for $K=2$, while leading to slightly larger or comparable error bars. This suggests that deep bound states described by an ab initio deduced NA optical potential can be reasonably well approximated by tri- to five-diagonal potentials, and in turn, can be successfully simulated on far-term error-corrected devices (practically yielding the true result) and even on NISQ processors (yielding energies within $1\sigma$ across a 14% $\hbar\omega$ variation).
$\hbar\omega$ $E_{0}$ $K$ $N$ $E_{\rm th}^{K}$ $E_{\rm NL}$ $E_{\rm shot}$ $E_{\rm noisy}$ $E_{\rm NR}$
@@ VI.5 Quantum simulations for n$+\alpha$ using ab initio optical potentials
$12$ $-18.85$ $1$ $8$ $-17.7986$ $-17.7986\pm 4\times 10^{-10}$ $-17.796\pm 0.001$ $31.6\pm 0.4$ $-16.8727$
$16$ $-17.7987$ $-17.7852\pm 0.0008$ $-17.791\pm 0.007$ $68.8\pm 0.7$ $-16.8073$
$2$ $8$ $-16.6190$ $-16.6189\pm 1.7\times 10^{-13}$ $-16.618\pm 0.001$ $58.7\pm 0.7$ $-15.8334$
$16$ $-16.6191$ $-16.615\pm 0.005$ $-16.60\pm 0.02$ $162\pm 2$ $-14.8717$
$16$ $-20.84$ $1$ $8$ $-20.7735$ $-20.7735\pm 3.2\times 10^{-11}$ $-20.7733\pm 0.0005$ $8.8\pm 0.3$ $-20.2597$
$2$ $8$ $-18.9470$ $-18.9470\pm 1.8\times 10^{-13}$ $-18.9470\pm 0.0006$ $43.2\pm 0.6$ $-18.3241$
@@
Figure 16: Simulation for the n-$\alpha$ ab initio potential for $\hbar\omega=12$ MeV Eq. (VI.2) with $N=16$, and (a) $K=1$ and (b) $K=2$. The inset plot is the last 200 iterations.
@@
Figure 17: Simulation for the n-$\alpha$ ab initio potential for $\hbar\omega=16$ MeV Eq. (VI.2) with $N=8$, and (a) $K=1$ and (b) $K=2$. The inset plot is the last 200 iterations.
@@ VII Conclusions and future directions
To estimate the measurement statistics, which is of key importance to successful simulations on quantum devices, we provide an extensive numerical analysis of the number of Pauli terms and qubit-wise commuting sets in the Hamiltonian as a function of the matrix size and the number of off-diagonals. We show that when the number of off-diagonals $12K+$ 1 exceeds the size of the matrix $N$, the number of Pauli terms and commuting sets saturate, with the Gray/binary encoding having fewer Pauli terms than the one-hot encoding. Beyond this point, more off-diagonals can be added to the problem, improving the approximation, without further load to the quantum device.
@@
To demonstrate the efficacy of the Gray encoding, the one-hot and Gray encodings are compared in quantum simulations of the lowest $\frac{1}{2}^{+}$ state of n+${}^{14}$, only bound by $1.04$ MeV in the $K=3$ approximate potential, for $N=8$ basis states. Indeed, compared to the Gray-code simulations with three qubits, the one-hot simulations with eight qubits are much slower and show worse performance in the presence of noise. In addition, we perform quantum simulations using a shot-noise and noisy simulator to inform the suitability of these simulations in the far-term error-corrected and NISQ regimes, respectively. It is remarkable that for n+${}^{10,12}$ and n+$\alpha$, we find that the shot-noise energies practically reproduce the corresponding exact value, whereas the noisy simulations coupled with the noise-resilient training method yield energies that deviate by less than one MeV. Finally, for the bound $\frac{1}{2}^{+}$ orbit of the neutron-alpha optical potential deduced ab initio, we report energy from the noisy quantum simulation that lies only within $1\sigma$ of the shot-noise outcome across a 14% $\hbar\omega$ variation.
@@
VIII Acknowledgements
SR thanks Prerna Agarwal, Stav Haldar, Kaiyuan Ji, Dhrumil Patel, Aby Philip, and Vishal Singh for helpful and illuminating discussions. SR specifically thanks Kaiyuan Ji for the idea behind the proof of Lemma A.8. This work was supported by the U.S. Department of Energy under award DE-SC0023694. We are grateful to M. Burrows for calculating the ab initio n+$\alpha$ non-local optical potentials, which were enabled by high performance computational resources provided by LSU, the National Energy Research Scientific Computing Center (under the DOE award DE-AC02-05CH11231), as well as the Frontera computing project (under the NSF award OAC-1818253). SR and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University, from the National Science Foundation under Grant No. 2315398, and from AFRL under agreement no. FA8750-23-2-0031.

### 15 Mar 2026 v3

@@ VI.3 Description of the quantum simulations
[…]
  • Exact diagonalization (referred to as “true value”): We diagonalize the matrix exactly (using classical methods) and find the minimum eigenvalue, which is used to validate the quantum simulation outcomes.
@@
  • Noise-resilient estimation: We use the final parameters $\bm{\theta}$ from the noisy simulation and calculate the expectation value of the Hamiltonian on a classical machine to find a noiseless estimate for the minimum energy, which we refer to as the noise-resilient (NR) value. The idea of noise-resilient training was first introduced in [44]. We find that the algorithm exhibits noise resilience; i.e., the NR value is more accurate than the noisy simulation outcome. In other words, training is still possible in a noisy scenario.
@@
[…]
@@ VI.4 Quantum simulations for neutron-Carbon dynamics
[…]
@@
$E_{\rm th}$ Encoding Figure NN $E_{{\rm th}}^{K=3}$ $E^{K=3}_{\rm NL}$ $E^{K=3}_{\rm shot}$ $E^{K=3}_{\rm noisy}$ $E^{K=3}_{\rm NR}$
@@
n+10C 6.78-6.78 Gray 11 88 6.5364-6.5364 $6.5364\pm 2\times 10^{-10}$ $6.5305\pm 0.0005$ $3.5\pm 0.5$ 5.9-5.9
$16$ 6.7346-6.7346 $668\pm 0.007 6.6$$62\pm 0.02$ $12.6\pm 0.7$ 6.0-6.0
n+12C 1.86-1.86 Gray 12 88 1.18495-1.18495 $184993\pm 1\times 10^{-6} 1.1$$1701\pm 0.0006$ $8.7\pm 0.5$ 0.5-0.5
$16$ 1.70020-1.70020 $1.6973\pm 0.0001$ $1.599\pm 0.003$ $13.9\pm 0.7$ 0.8-0.8
n+14C 1.22-1.22 Gray 13 88 0.49963-0.49963 $0.4996\pm 6\times 10^{-10}$ $0.4966\pm 0.0005$ $8.4\pm 0.3$ $0.1$
$16$ 1.0070-1.0070 $0.9860\pm 0.0005$ $0.982\pm 0.002$ $15.2\pm 0.5$ $0.6$
One-hot 14 88 0.49963-0.49963 $0.4987\pm 0.0001$ $0.49\pm 0.03$ $9.0\pm 0.6$ $0.7$
Table 4: Simulation energy for the lowest $\frac{1}{2}^{+}$ state for each neutron-Carbon system, for the $K=3$ approximation with $N=8$ or $N=16$, corresponding to Figs. 11-14: the true value $E_{\rm th}^{K=3}$; $E^{K=3}_{\rm NL}$ from the noiseless simulations, $E^{K=3}_{\rm shot}$ from the shot-noise simulations, and $E^{K=3}_{\rm noisy}$ from the noisy simulations, reported as the mean and standard deviation of the last 100 iterations; and the NR value $E^{K=3}_{\rm NR}$. The theoretical eigenenergy $E_{\rm th}$ of the exact potential is also shown (cf. Table 3).
In particular, in all cases the noiseless and shot-noise simulations with the Gray code yield results with very small errors, mostly, $10^{-10}$-$10^{-3}$. Importantly, they reproduce the true value within 0-22%, with the only exception being the 66% deviation in $E^{K=3}_{\rm shot}$ for 12C and $N=16$.
@@
We note here that for the example of n+10C, with $N=8$ and $K=3$, in addition to the noise resilient final estimate, we report the mean and standard deviation of the noiseless estimates of the last 100 noisy iterations. In other words, we use the parameters from last 100 noisy iterations, calculate the noiseless estimate, and report the mean and standard deviation to be $6.05\pm 0.07$. This is $\sim 2\sigma$ away from the $E^{K=3}_{NR}$ value reported in Table 4 calculated from the last iteration only, which is very reasonable. This further solidifies the idea that the noise resilient method offers a much better estimate as compared to the noisy estimate. However, in any large-scale experiment, we would not recommend running the noiseless simulation for hundreds of iterations. This is because each noiseless simulation is prohibitively expensive. Instead, in this study, we use the standard deviation for n+10C as a guidance to the number of the significant digits we report for the noise resilient estimates in all simulations.
Furthermore, even in the case of weakly bound states, the Gray code with three qubits is superior to the one-hot encoding with eight qubits, as illustrated in Fig. 14 and Table 4. Indeed, compared to the Gray-code case, the one-hot simulations are much slower and show worse performance in the presence of noise. In particular, the one-hot simulation yields larger errors [e.g., by six (two) orders of magnitude for the noiseless (shot noise) simulations], as well as a deviation for $E^{K=3}_{\rm NR}$ from the true value that is twice as large as the corresponding Gray-code estimate. In addition, as shown in Figs. 2 and 3 as well as in Tables 1 and 2, the one-hot ansatz utilizes 13 two-qubit gates compared to only 8 two-qubit (CNOT) gates in the Gray encoding ansatz with $L=4$ used in our simulations. This suggests that the Gray encoding and the use of fewer qubits are indeed highly advantageous for nuclear problems that achieve convergence in larger model spaces, with the case of weakly bound systems presented here being an illustrative example.
@@
Figure 15: Simulation with the Gray encoding for the n-$\alpha$ ab initio potential for $\hbar\omega=12$ MeV Eq. (VI.2) with $N=8$ ($n=3$), and (a) $K=1$ and (b) $K=2$. The different types of simulations are detailed in Sec. VI.3. The inset plots the last 200 iterations.
[…]
We summarize the energy estimates from the various quantum simulations in Table 5. The energies $E_{\rm th}^{K}$ from the exact diagonalization for $K=1$ and $2$ converge within the first four to five digits with the increasing model space from $N=8$ to $N=16$, as illustrated in Table 5 for $\hbar\omega=12$ MeV (for $\hbar\omega=16$ MeV, $E_{\rm th}^{K=1}=-20.77$ MeV and $E_{\rm th}^{K=2}=-18.95$ MeV for $N=16$). Hence, the final estimates are based on the $N=8$ results, where the uncertainties are estimated for a 14% variation of the $\hbar\omega$ values ($\hbar\omega=12$-$16$ MeV), the same range that yields a total cross section for the neutron scattering on 4He that is converged and in agreement with experiment, as discussed in Ref. [11]. In addition, the $K=1$ and $K=2$ approximation energies are very close to the corresponding $E_{0}$ eigenenergy of the original non-local ab initio optical potential. In fact, across the $\hbar\omega$ range, the energy of the lowest $\frac{1}{2}^{+}$ orbit is estimated at 19.3(1.5)-19.3(1.5) MeV for $K=1$ and at 17.8(1.2)-17.8(1.2) MeV for $K=2$, both of which agree within the uncertainties with the non-local estimate of 19.8(1.0)-19.8(1.0) reported above. Exactly the same energy estimates are obtained by the noiseless and shot simulations ($19.3\pm 1.5$ MeV for $K=1$ and at $17.8\pm 1.2$ MeV for $K=2$). We note that the errors that stem from these simulations (of order of a few keV) are inconsequential compared to the ones associated with the $\hbar\omega$ variation ($\sim 1-1.5$ MeV). Importantly, while the noisy simulations yield unacceptably large energies, the corresponding NR energies agree with the shot-noise outcomes within $1\sigma$: 18.6(1.7)-18.6(1.7) MeV for $K=1$ and 17.1(1.2)-17.1(1.2) MeV for $K=2$, while leading to slightly larger or comparable error bars. This suggests that deep bound states described by an ab initio deduced NA optical potential can be reasonably well approximated by tri- to five-diagonal potentials, and in turn, can be successfully simulated on far-term error-corrected devices (practically yielding the true result) and even on NISQ processors (yielding energies within $1\sigma$ across a 14% $\hbar\omega$ variation).
$\hbar\omega$ $E_{0}$ NN Figure $K$ $E_{\rm th}^{K}$ $E_{\rm NL}$ $E_{\rm shot}$ $E_{\rm noisy}$ $E_{\rm NR}$
[MeV] [MeV] [MeV] [MeV] [MeV] [MeV]
$12$ 18.85-18.85 88 15 11 17.7986-17.7986 $17.7986\pm 4\times 10^{-10}$ $17.796\pm 0.001$ $31.6\pm 0.4$ 16.9-16.9
22 16.6190-16.6190 $6189\pm 1.7\times 10^{-13} 16.61$$618\pm 0.001$ $58.7\pm 0.7$ 15.8-15.8
$16$ 16 11 17.7987-17.7987 $17.7852\pm 0.0008$ $17.791\pm 0.007$ $68.8\pm 0.7$ 16.8-16.8
22 16.6191-16.6191 $16.615\pm 0.005$ $16.60\pm 0.02$ $162\pm 2$ 14.9-14.9
$16$ 20.84-20.84 88 17 11 20.7735-20.7735 $20.7735\pm 3.2\times 10^{-11}$ $20.7733\pm 0.0005$ $8.8\pm 0.3$ 20.3-20.3
22 18.9470-18.9470 $18.9470\pm 1.8\times 10^{-13}$ $18.9470\pm 0.0006$ $43.2\pm 0.6$ 18.3-18.3
[…]
Figure 16: Simulation with the Gray encoding for the n-$\alpha$ ab initio potential for $\hbar\omega=12$ MeV Eq. (VI.2) with $N=16$ ($n=4$), and (a) $K=1$ and (b) $K=2$. The different types of simulations are detailed in Sec. VI.3. The inset plots the last 200 iterations.
[…]
Figure 17: Simulation for the n-$\alpha$ ab initio potential for $\hbar\omega=16$ MeV Eq. (VI.2) with $N=8$, and (a) $K=1$ and (b) $K=2$. The different types of simulations are detailed in Sec. VI.3. The inset plots the last 200 iterations.
VI.6 Comparing QC and DGC schemes
In this section we compare the efficacy of QC and DGC schemes using both the Gray and binary encodings. We perform both noiseless and shot noise simulations for all cases. In the shot noise runs, we keep the total number of shots per variational run to be fixed. This is to simulate a fixed amount of “quantum” resources. Furthermore, in contrast with the preceding simulations in this work, here we only use gradient descent with a varying learning rate scheme to estimate the gradient. This removes the effects of the SPSA noise and gives a clearer picture of the inherent shot noise. However, this also leads to results of the noiseless simulations in this section being slightly different from those reported in Table 5, since the noise introduced by SPSA can take the system out of a local minima and converge to a better solution.
[…]
Simulation Mean and Standard Deviation
Noiseless + QC $17.78384\pm 0.00009$
Noiseless + DGC $17.78384\pm 0.00009$
Shot Noise + QC $17.67\pm 0.14$
Shot Noise + DGC $17.72\pm 0.13$
[…]
Next, we compare the Gray and binary encoding using the QC scheme. From the results of Sec. V.2, we expect the Gray encoding to outperform the binary encoding, and the data in Fig. 19 clearly corroborates this. The outcomes of these simulations are summarized in Table 7. While the Gray encoding agrees with the noiseless result within $1\sigma$, the binary encoding leads to a larger standard deviation and at least a $2\sigma$ agreement.
Simulation Mean and Standard Deviation
Noiseless $16.5990\pm 8\times 10^{5}$
Shot Noise + QC + Gray $16.31\pm 0.34$
Shot Noise + QC + binary $15.80\pm 0.57$
[…]
@@ VII Conclusions and future directions
To estimate the measurement statistics, which is of key importance to successful simulations on quantum devices, we provide an extensive numerical analysis of the number of Pauli terms and qubit-wise commuting sets in the Hamiltonian as a function of the matrix size and the number of off-diagonals. We show that when the number of off-diagonals $2K+1$ exceeds the size of the matrix $N$, the number of Pauli terms and commuting sets saturate, with the Gray/binary encoding having fewer Pauli terms than the one-hot encoding. Beyond this point, more off-diagonals can be added to the problem, improving the approximation, without further load to the quantum device.
@@
[…]
@@
VIII Code Availability
The code for the simulations can be found as a Zenodo repository [40].
IX Acknowledgements
SR thanks Prerna Agarwal, Stav Haldar, Kaiyuan Ji, Dhrumil Patel, Aby Philip, and Vishal Singh for helpful and illuminating discussions. SR specifically thanks Kaiyuan Ji for the idea behind the proof of Lemma A.8. This work was supported by the U.S. Department of Energy under award DE-SC0023694. We are grateful to M. Burrows for calculating the ab initio $\alpha n$on-local optical potentials, which were enabled by high performance computational resources provided by LSU, the National Energy Research Scientific Computing Center (under the DOE award DE-AC02-05CH11231), as well as the Frontera computing project (under the NSF award OAC-1818253). SR and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University, from the National Science Foundation under Grant No. 2315398, and from AFRL under agreement no. FA8750-23-2-0031.